OpenStudy (prince1342):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

I guess slope-intercept form is y = kx + m for some constants k, m, because k is slope and m is y-intercept.

OpenStudy (anonymous):

You need to get rid of everything that has to do with x on the left-hand side, so subtract 2x from both sides to get a new equation which is equivalent to the original.

OpenStudy (anonymous):

You then have only 3y at the left-hand side. Get rid of the coefficient of y by dividing both sides by 3. This gives you the final equation y = -2x/3 + 400.

OpenStudy (texaschic101):

actually, slope intercept is y = mx + b, and m is your slope and b is your y intercept. But you are doing it correctly and I will let you finish :)

OpenStudy (anonymous):

As for showing all of your work, I like to put what I do next on the right of equations, like so: \[2x + 3y = 1200 \text{ [subtract 2x]}\\ 3y = -2x + 1200 \text{ [divide by 3]}\\ y = -\frac{2}{3}x + 400\]